OpenStudy (anonymous):

Math Check please :)! \[6x ^{4}-11x ^{2}+3=0\] I got 1/3,3/2

jimthompson5910 (jim_thompson5910):

did you let z = x^2 (or something like that)? If so, then you found the solutions for z, but not x

OpenStudy (anonymous):

Yeah, looks like that's what happened :) ^^^

OpenStudy (anonymous):

you are right, x = (1/3) and (3/2)

jimthompson5910 (jim_thompson5910):

so you need to solve x^2 = 1/3 and x^2 = 3/2 Note: you should get 4 solutions

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

thats what I did :D! thanks haha but now the question is, how do you use that to find x?

OpenStudy (anonymous):

oh thanks Jim! tons of help!

jimthompson5910 (jim_thompson5910):

see my second post

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

yeah kdsim.. when you have 4 degree there will be 4 roots